# Double Bangs at IceCube as a Window to the Neutrino Mass Origin

### v1

@@ Abstract
In this Letter, we will specifically show that RG effects can enhance the high-energy tau neutrino flux at neutrino telescopes compared to the expectation in the absence of new physics. We will consider both atmospheric and astrophysical neutrinos, where the appearance of tau neutrinos from the former arises exclusively from zero-baseline flavor transitions, characteristic of scenarios with RG running as discussed above. An excess of tau neutrinos implies that more “double bang” events are expected to be recorded in the detector. These events consist of two spatially separated showers, typically displaced by $\sim 50\text{ m/PeV}$: the first originates from a tau neutrino charged-current interaction that produces a tau lepton, while the second arises from the subsequent decay of the tau lepton after propagating a macroscopic distance. We will present a statistically significant excess of new physics-induced double bangs for a number of benchmark scenarios from a variant of the scotogentic model of neutrino mass introduced in Ref. Babu:2021cxe . Before turning to the details of this analysis, let us note the complementarity of our approach with Refs. Coloma:2017ppo ; Coloma:2019qqj ; Atkinson:2021rnp , where new physics-induced double bangs are generated by the production of heavy right-handed neutrinos that decay after propagating macroscopic distances. In our scenario, new physics particles are not directly produced but enter only through loop diagrams that yield the RG running of the PMNS matrix.
@@
Now, if we flip the script and consider $N_{R}$ and $\varphi$ with masses equal to $Q_{\text{NuTeV}}$, $Y_{N}$ would be subject to RG running only above $Q_{\text{NuTeV}}$. In such a setup, where we effectively have $U(Q_{\text{NuTeV}}^{2})=U(Q_{p}^{2})$ and the mismatch takes place when $Q>Q_{\text{NuTeV}}$, NuTeV and other short-baseline experiments do not impose any constraints since nonvanishing zero-baseline oscillation probabilities are only generated above their typical momentum transfer scales for neutrino scattering. At such values of $Q_{d}$, we will focus on high-energy atmospheric neutrinos and astrophysical neutrinos at neutrino telescopes. The zero-baseline discussion applies to the former. For definiteness, we will focus on IceCube, since its flavor efficiencies are publicly available for high-energy events. A relevant benchmark point for the IceCube High-Energy Starting Events (HESE) IceCube:2020wum selection is shown in Fig. 1 with solid lines. All the benchmarks are tabulated in the Supplemental Material for the purpose of reproducibility of these results. We observe that $P_{\alpha\beta}$ (solid) can reach $\mathcal{O}(0.1)$ without violating the NuTeV constraints, and such values are large enough to record an observable impact of the running, as we will demonstrate in the remainder of this work.
@@
where $T$ is the data-taking time, $\epsilon_{\tau}(E_{\nu})$ is the detector efficiency for identifying $\nu_{\tau}$, $\Phi_{\alpha}$ is the neutrino flux (either atmospheric or astrophysical), and $\sigma_{\nu N}$ is the neutrino-nucleus deep inelastic scattering (DIS) cross section. Further, $w_{\alpha\to\tau}^{\mathrm{BSM}}$ is what we denote as “BSM weight,” and it is closely related to the neutrino flavor transition probability. In fact, if Eqs. 8 and 7 were not $Q^{2}$-dependent, this factor would have been exactly equal to the respective transition probabilities. However, given the momentum transfer dependence in $P_{\alpha\tau}$, the transition probabilities need to be weighted with the $Q^{2}$-dependent differential cross section $\frac{d^{2}\sigma_{\nu(\bar{\nu})}}{dx\,dQ^{2}}$  Formaggio:2012cpf ; Xie:2023suk , and this is precisely how the BSM weight is computed. The isoscalar $u(\bar{u})$ and $d(\bar{d})$ PDFs, that appear in the differential cross section, are calculated from CT18 NNLO proton PDFs Buckley:2014ana . For details of the BSM weight, we refer the reader to the Supplemental Material.
@@
where $X_{\alpha}^{\rm prod}$ represents the flavor fraction of a given neutrino type at the production site, and $\Phi^{\rm astro}_{1\nu}(E_{\nu})=\frac{1}{6}\Phi^{\rm astro}_{6\nu}(E_{\nu})$, is the (anti)neutrino flux for a single flavor, obtained from the single power law formula from (IceCube:2020wum, ).
For a given parameter point, including both atmospheric and astrophysical neutrino fluxes, we can compare the expected number of double bangs in the presence of RG running with the result in the case with no RG effects; the latter is computed by setting BSM weights $w_{\alpha\to\tau}^{\mathrm{BSM}}$ to $\delta_{\alpha\tau}$}. We did this for $\mathcal{O}(10^{3})$ parameter points in the model, and the results are shown in Fig. 2, displaying the difference in number of events between the running and the standard case, i.e., $\Delta N_{\alpha}=N^{\rm BSM}_{\alpha}-N^{\rm std}_{\alpha}$. All parameter points, event differences, and the results for different initial flavor compositions of astrophysical neutrinos are available on github . We found that the results are only mildly dependent on the flavor ratio at the production site; we present $\nu_{e}:\nu_{\mu}:\nu_{\tau}=1:2:0$ in Fig. 2. On the y-axis, we show an enhancement of double bangs with respect to the standard expectation, $\Delta N_{\tau}$. On the x-axis, for illustration, we show the difference in $\nu_{\mu}$-like track events for the corresponding parameter points, computed analogously to tau events (see Eqs. 10 and 11), only with different weights to account for different transition probabilities and detection efficiencies. The $\Delta N_{\mu}$ values reported in Fig. 2 are consistent with the measurement of the atmospheric muon neutrino flux at IceCube IceCube:2014slq . We observe a number of parameter points for which RG running yields $\sim 2$ additional double bangs, with a single parameter point reaching $\Delta N_{\tau}=2.5$ (labeled with a star). For this benchmark point, we show the RG evolution for mixing angles and phases as well as the reconstructed energy spectrum of tau-like events in the Supplemental Material.
@@ 1. BSM weights
At the considered energies, the neutrino-nucleus interaction is governed by deep inelastic scattering (DIS). A neutrino $\nu_{\alpha}$ can produce $\ell_{\alpha}$ through the charged-current interaction in which a $W$ boson is exchanged. The flavor-independent charged-current DIS differential cross section reads
$\frac{d^{2}\sigma_{\nu(\bar{\nu})}}{dx\,dQ^{2}}=\frac{G_{F}^{2}}{4\pi x\left(1+Q^{2}/M_{W}^{2}\right)^{2}}\left[Y_{+}F_{2}-y^{2}F_{L}\pm Y_{-}xF_{3}\right],$ (A1)
where $F_{2,3,L}$ are structure functions, $Y_{\pm}$ are functions of kinematic invariants, and $x$ and $y={Q^{2}}/{(2xm_{N}E_{\nu})}$ are Bjorken variables.
We employ the DIS cross section (up to leading order, $F_{L}=0$) to compute the BSM weights $w_{\alpha\to\beta}^{\mathrm{BSM}}(E_{\nu})$, in order to properly treat the RG evolution–modified flavor transitions. Specifically, we define the BSM weight as the ratio of the integrated DIS differential cross section weighted by the neutrino transition probability to the standard DIS cross section
$w_{\alpha\to\beta}^{\mathrm{BSM}}(E_{\nu})=\frac{\int_{\nu}\,P_{\alpha\beta}(E_{\nu},Q^{2})+\int_{\bar{\nu}}\,P_{\bar{\alpha}\bar{\beta}}(E_{\nu},Q^{2})}{\int_{\nu}\quad+\quad\int_{\bar{\nu}}}\,$, (A2)
with
$\int_{\nu(\bar{\nu})}=\int_{Q_{\text{min}}^{2}}^{Q_{\text{max}}^{2}}\,dQ^{2}\,\int_{Q^{2}/Q_{\text{max}}^{2}}^{1}\,dx\,\frac{d^{2}\sigma_{\nu(\bar{\nu})}}{dx\,dQ^{2}}\,.$ (A3)
Here, $Q_{\text{max}}^{2}=2m_{N}E_{\nu}$ and $Q_{\text{min}}=1.3~\text{GeV}$ (from CT18 NNLO).
@@ 2. Benchmarks
$Y_{\nu}=\frac{1}{\sqrt{C}}\,U^{\dagger}(Q_{p}^{2})\,\sqrt{\text{diag}\left(m_{\nu_{1}},m_{\nu_{2}},m_{\nu_{3}}\right)}\,R\,Y_{N}^{-1/2}\,$, (A4)
where $C$ is the VEV-dependent prefactor in Eq. 3. The orthogonal matrix $R$ is computed by three Euler-like rotations, $R=R_{12}(\xi_{1})R_{23}(\xi_{2})R_{13}(\xi_{3})$ with rotation angles $\xi_{1}$ in the 1-2 plane, $\xi_{2}$ in the 2-3 plane, and $\xi_{3}$ in the 1-3 plane. In addition to $\xi_{i}$ and $Y_{N}$, the lightest neutrino mass enters as a free parameter in the diagonal mass matrix in Eq. A4.
We present the model parameters (rounded to three significant digits) at the production scale. They correspond to three benchmark points. Two are used in Fig. 1 (dotted and solid lines) and are labeled here as $Q_{\pi}$’ and $Q_{\text{NuTeV}}$’, respectively. The third one is the $\bigstar$ benchmark from Fig. 2.
@@
SUPPL. TABLE 1: Leptonic mixing parameters and neutrino mass-squared differences for the selected benchmark points.
@@
SUPPL. TABLE 2: Rotation angles in the Casas-Ibarra parametrization, the Yukawa couplings, and lightest neutrino mass for the selected benchmark points. We consider the normal mass ordering.
3. RG running and reconstructed spectrum for $\bigstar$} benchmark point
Refer to caption Refer to caption
SUPPL. FIG. 1: In the left panel, we show the RG evolution of mixing angles and phases as defined in Eq. 5. The angle $\theta_{12}$, along with all three phases, changes significantly in the relevant $Q^{2}$ range, while the angles $\theta_{13}$ and $\theta_{23}$ do not, in accord with the arguments in Ref. Antusch:2005gp . In the right panel, we show the corresponding reconstructed energy spectrum of the tau-like events. The atmospheric contribution from $\nu_{e}$ is too small to be observed in the presented range. If one subtracts the unfilled (no running) histogram from the stacked one and integrates over energies, $\Delta N_{\tau}=2.5$ from Fig. 2 is obtained.

### v3

@@ Abstract
In this Letter, we show that RG effects can increase the high-energy tau neutrino flux at neutrino telescopes compared to the prediction without new physics. We demonstrate this using the variant of the scotogenic neutrino mass model introduced in Ref. Babu:2021cxe. We consider both atmospheric and astrophysical neutrinos, where the appearance of tau neutrinos from the former arises exclusively from zero-baseline flavor transitions, characteristic of scenarios with RG running as discussed above. An excess of tau neutrinos implies that more “double bang” events are expected to be recorded in the detector. These events consist of two spatially separated showers, typically displaced by $\sim 50\text{ m/PeV}$: the first originates from a tau neutrino charged-current interaction that produces a tau lepton, while the second arises from the subsequent decay of the tau lepton after propagating a macroscopic distance. Before turning to the details of our analysis, let us note the complementarity of our approach with Refs. Coloma:2017ppo; Coloma:2019qqj; Atkinson:2021rnp, where new physics-induced double bangs are generated by the production of heavy right-handed neutrinos that decay after propagating macroscopic distances. In our scenario, new physics particles are not directly produced but enter only through loop diagrams that yield the RG running of the PMNS matrix.
@@
Now, if we flip the script and consider $N_{R}$ and $\varphi$ with masses equal to $Q_{\text{NuTeV}}$, $Y_{N}$ would be subject to RG running only above $Q_{\text{NuTeV}}$. In such a setup, where we effectively have $U(Q_{\text{NuTeV}}^{2})=U(Q_{p}^{2})$ and the mismatch takes place when $Q>Q_{\text{NuTeV}}$, NuTeV and other short-baseline experiments do not impose any constraints since nonvanishing zero-baseline oscillation probabilities are only generated above their typical momentum transfer scales for neutrino scattering. At such values of $Q_{d}$, we will focus on high-energy atmospheric neutrinos and astrophysical neutrinos at neutrino telescopes. The zero-baseline discussion applies to the former. For definiteness, we will focus on IceCube, since its flavor efficiencies are publicly available for high-energy events. A relevant benchmark point for the IceCube High-Energy Starting Events (HESE) IceCube:2020wum selection is shown in Fig. 1 with solid lines. All the benchmarks are tabulated in the appendix for the purpose of reproducibility of these results. We observe that $P_{\alpha\beta}$ (solid) can reach $\mathcal{O}(0.1)$ without violating the NuTeV constraints, and such values are large enough to record an observable impact of the running, as we will demonstrate in the remainder of this work.
@@
where $T$ is the data-taking time, $\epsilon_{\tau}(E_{\nu})$ is the detector efficiency for identifying $\nu_{\tau}$, $\Phi_{\alpha}$ is the neutrino flux (either atmospheric or astrophysical), and $\sigma_{\nu N}$ is the neutrino-nucleus deep inelastic scattering (DIS) cross section. Further, $w_{\alpha\to\tau}^{\mathrm{BSM}}$ is what we denote as “BSM weight,” and it is closely related to the neutrino flavor transition probability. In fact, if Eqs. 8 and 7 were not $Q^{2}$-dependent, this factor would have been exactly equal to the respective transition probabilities. However, given the momentum transfer dependence in $P_{\alpha\tau}$, the transition probabilities need to be weighted with the $Q^{2}$-dependent differential cross section $\frac{d^{2}\sigma_{\nu(\bar{\nu})}}{dx\,dQ^{2}}$  Formaggio:2012cpf; Xie:2023suk, and this is precisely how the BSM weight is computed. The isoscalar $u(\bar{u})$ and $d(\bar{d})$ PDFs, that appear in the differential cross section, are calculated from CT18 NNLO proton PDFs Hou:2019efy. For details of the BSM weight, we refer the reader to the appendix.
Figure 3: In the left panel, we show the RG evolution of mixing angles and phases as defined in Eq. 5. In the right panel, we show the corresponding reconstructed energy spectrum of the tau-like events.
@@
where $X_{\alpha}^{\rm prod}$ represents the flavor fraction of a given neutrino type at the production site, and $\Phi^{\rm astro}_{1\nu}(E_{\nu})=\frac{1}{6}\Phi^{\rm astro}_{6\nu}(E_{\nu})$, is the (anti)neutrino flux for a single flavor, obtained from the single power law formula from (IceCube:2020wum).
For a given parameter point, including both atmospheric and astrophysical neutrino fluxes, we can compare the expected number of double bangs in the presence of RG running with the result in the case with no RG effects; the latter is computed by setting BSM weights $w_{\alpha\to\tau}^{\mathrm{BSM}}$ to $\delta_{\alpha\tau}$}. We did this for $\mathcal{O}(10^{3})$ parameter points in the model, and the results are shown in Fig. 2, displaying the difference in number of events between the running and the standard case, i.e., $\Delta N_{\alpha}=N^{\rm BSM}_{\alpha}-N^{\rm std}_{\alpha}$. All parameter points, event differences, and the results for different initial flavor compositions of astrophysical neutrinos are available on github  . We found that the results are only mildly dependent on the flavor ratio at the production site; we present $\nu_{e}:\nu_{\mu}:\nu_{\tau}=1:2:0$ in Fig. 2. On the y-axis, we show an enhancement of double bangs with respect to the standard expectation, $\Delta N_{\tau}$. On the x-axis, for illustration, we show the difference in $\nu_{\mu}$-like track events for the corresponding parameter points, computed analogously to tau events (see Eqs. 10 and 11), only with different weights to account for different transition probabilities and detection efficiencies. The $\Delta N_{\mu}$ values reported in Fig. 2 are consistent with the measurement of the atmospheric muon neutrino flux at IceCube IceCube:2014slq. We observe a number of parameter points for which RG running yields $\sim 2$ additional double bangs, with a single parameter point reaching $\Delta N_{\tau}=2.5$ (labeled with a star).
For this benchmark point, we show the RG evolution for mixing angles and phases as well as the reconstructed energy spectrum of tau-like events in Fig. 3. The angle $\theta_{12}$, along with all three phases, changes significantly in the relevant $Q^{2}$ range in the left panel of Fig. 3, while the angles $\theta_{13}$ and $\theta_{23}$ do not, in accord with the arguments in Ref. Antusch:2005gp. The right panel compares the case with no running to the case where additional double bang events arise from atmospheric and astrophysical neutrinos in the presence of RG effects. The atmospheric contribution from $\nu_{e}$ is too small to be observed in the presented range. If one subtracts the unfilled (no running) histogram from the stacked one and integrates over energies, $\Delta N_{\tau}=2.5$ from Fig. 2 is obtained.
@@ 1. BSM weights
At the considered energies, the neutrino-nucleus interaction is governed by deep inelastic scattering (DIS). A neutrino $\nu_{\alpha}$ can produce $\ell_{\alpha}$ through the charged-current interaction in which a $W$ boson is exchanged. The flavor-independent charged-current DIS differential cross section reads Formaggio:2012cpf; Xie:2023suk
$\frac{d^{2}\sigma_{\nu(\bar{\nu})}}{dx\,dQ^{2}}=\frac{G_{F}^{2}}{4\pi x\left(1+Q^{2}/M_{W}^{2}\right)^{2}}\left[Y_{+}F_{2}-y^{2}F_{L}\pm Y_{-}xF_{3}\right],$ (12)
where $F_{2,3,L}$ are structure functions, $Y_{\pm}$ are functions of kinematic invariants, and $x$ and $y={Q^{2}}/{(2xm_{N}E_{\nu})}$ are Bjorken variables. We employ the DIS cross section (up to leading order, $F_{L}=0$) to compute the BSM weights $w_{\alpha\to\beta}^{\mathrm{BSM}}(E_{\nu})$, in order to properly treat the RG evolution–modified flavor transitions. Specifically, we define the BSM weight as the ratio of the integrated DIS differential cross section weighted by the neutrino transition probability to the standard DIS cross section
$w_{\alpha\to\beta}^{\mathrm{BSM}}(E_{\nu})=\frac{\int_{\nu}\,P_{\alpha\beta}(E_{\nu},Q^{2})+\int_{\bar{\nu}}\,P_{\bar{\alpha}\bar{\beta}}(E_{\nu},Q^{2})}{\int_{\nu}\quad+\quad\int_{\bar{\nu}}}\,.$ (13)
The integration $\int_{\nu(\bar{\nu})}$ evaluated with $Q_{\text{max}}^{2}=2m_{N}E_{\nu}$ and $Q_{\text{min}}=1.3~\text{GeV}$ (CT18 NNLO) follows
$\int_{\nu(\bar{\nu})}=\int_{Q_{\text{min}}^{2}}^{Q_{\text{max}}^{2}}\,dQ^{2}\,\int_{Q^{2}/Q_{\text{max}}^{2}}^{1}\,dx\,\frac{d^{2}\sigma_{\nu(\bar{\nu})}}{dx\,dQ^{2}}\,.$ (14)
@@ 2. Benchmarks
$Y_{\nu}=\frac{1}{\sqrt{C}}\,U^{\dagger}(Q_{p}^{2})\,\sqrt{\text{diag}\left(m_{\nu_{1}},m_{\nu_{2}},m_{\nu_{3}}\right)}\,R\,Y_{N}^{-1/2}\,$, (15)
where $C$ is the VEV-dependent prefactor in Eq. 3. The orthogonal matrix $R$ is computed by three Euler-like rotations, $R=R_{12}(\xi_{1})R_{23}(\xi_{2})R_{13}(\xi_{3})$ with rotation angles $\xi_{1}$ in the 1-2 plane, $\xi_{2}$ in the 2-3 plane, and $\xi_{3}$ in the 1-3 plane. In addition to $\xi_{i}$ and $Y_{N}$, the lightest neutrino mass enters as a free parameter in the diagonal mass matrix in Eq. 15. We present the model parameters (rounded to three significant digits) at the production scale. They correspond to three benchmark points. Two are used in Fig. 1 (dotted and solid lines) and are labeled here as $Q_{\pi}$’ and $Q_{\text{NuTeV}}$’, respectively. The third one is the $\bigstar$ benchmark from Fig. 2.
@@
Table 1: Leptonic mixing parameters and neutrino mass-squared differences for the selected benchmark points.
@@
Table 2: Rotation angles in the Casas-Ibarra parametrization, the Yukawa couplings, and lightest neutrino mass for the selected benchmark points. We consider the normal mass ordering.